OpenStudy (anonymous):

Choose the correct simplification of (3x3 + 2x2 − 5x) − (8x3 − 2x2)

OpenStudy (anonymous):

A) −5x3 + 4x2 − 5x B) 5x3 + 2x2 + 5x C) −5x3 − 4x2 D) 5x3 + 2x2

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 3x^3 + 2x^2 - 5x - 8x^3 + 2x^2 }\) \(\Large \color{purple}{\rightarrow -5x^3 + 4x^2 - 5x }\)

OpenStudy (anonymous):

I HAVE ANOTHER

Parth (parthkohli):

Always remember that when the brackets get removed and a negative is outside the bracket, then the signs get changed in the bracket.

OpenStudy (diyadiya):

Post your question !~

OpenStudy (apoorvk):

Add/subtract coefficients of terms with alike degrees of 'x'. For example, 8x^3 -5x^3 = (8-5)x^3 = 3x^3. Same for the others. Think you can do this now?

OpenStudy (anonymous):

Diyadiya!@!@

OpenStudy (anonymous):

@Diyadiya